OpenStudy (anonymous):

15x2 - 44x + 21 = 0 A. x = B. x = 5 C. x = D. x = E. x = 3 F. x =

OpenStudy (lgbasallote):

rewrite...

OpenStudy (anonymous):

G. Don't hate math... :D

OpenStudy (anonymous):

x= 3/5, x =7/3

OpenStudy (anonymous):

Have you learned the quadratic equation?